OpenStudy (anonymous):

solve by the subtitution method 5x+7y=15 x=26-6y

OpenStudy (angela210793):

5(26-6y)+7y=15 130-23y=15 -23y=-115 y=5 x=26-30=-4

OpenStudy (angela210793):

solution (-4;5)

OpenStudy (anonymous):

5x+7y=15 =>5(26-6y)+7y=15 =>130-30y+7y=15 =>23y=115 =>y=5 =>x=-4